OpenStudy (anonymous):

solve and show work: 12 is what % of 60

OpenStudy (anonymous):

let it be x then x % of 60=12 (x/100) * 60=12 solve for x

OpenStudy (anonymous):

ie x/100=12/60 x/100=1/5 x=100/5 x=20 20 % ans.. hope you get it @purplecows

OpenStudy (anonymous):

The answer is 20. Formula: is over of ? over 100